Use simpson's method to approximate integral 0 to 8(x^2-x+3)dx with 8 such-intervals.
1
Expert's answer
2020-05-18T18:50:53-0400

"simpson's" "rule" ,

"y = x^{2}-x+3"

"integral"

"a =0"

"b =8"

"interval"

"n =8"

"Area = \int_{a}^{b} f(x)dx"


"=\frac{\triangle x}{3} (y_{0}+4y_{1}+2y_{2}+4y_{3}+2y_{4}....+4y_{n-1}+y_{n})"


"where \triangle x = \frac{b-a}{n}"


"y_{0} = 0-0+3 =3"

"y_{1} = 3"

"y_{2} = 5"

"y_{3} = 9"

"y_{4} = 15"

"y_{5} = 23"

"y_{6} = 33"

"y_{7} = 45"

"y_{8} = 59"

"\triangle x = \frac{8-0}{8}"

"\triangle x =1"

"area =\frac{1}{3} (3+4\times3+2\times5+4\times 9+2\times15+4\times23+2\times33+4\times45+59)"


 "area = 162.67"